Circular No. 5683

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1992bq IN ANONYMOUS GALAXY M. Hamuy, Cerro Tololo Interamerican Observatory; and J. Maza,

University of Chile, report the discovery of an apparent supernova (m_{pg} ∼ 17) by R. Antezana (University of Chile) on a 20-min unfiltered IIa-O plate taken by G. Valladares with the CTIO Curtis Schmidt telescope on Dec. 20.32 UT. SN 1992bq is located 15′′ east and 1′′ south of the nucleus of a spiral galaxy at α = 10^{h}14^{m}43^{s}.6, δ = −34^{o}30′08′′ (equinox 1950.0). Confirming observations were made by R. Aviles from CCD B, V, R, and I images obtained on Dec. 23.31 with the CTIO 0.9-m telescope.

NOVA CYGNI 1992 C. E. Woodward, University of Wyoming; and M. A. Greenhouse, Na-

tional Air and Space Museum, Smithsonian Institution, write: “Analysis of infrared spectra of N Cyg 1992 obtained on Sept. 8.3 UT using the 2.1-m telescope (+ CRSP) at Kitt Peak National Observatory (IAUC 5612) have revealed an emission line observed at 1.250 ± 0.002 μm. We tentatively identify this feature as the 2s^{2}2p^{4} ^{3}P_{2}-^{3}P_{1} transition of [S IX]. Strong hydrogen Paβ emission also was present. Astronomical observation of the [S IX] transition has not been reported previously. This line may be accompanied by the ^{3}P_{1}-^{3}P_{0} transition of [S IX] at 3.75 μm, since its excitation temperature is only 4000 K higher than that of the ^{3}P_{2}-^{3}P_{1} transition. However, the critical density for collisional de-excitation of the 3.75-μm line is approximately a factor of 4 lower than that of the 1.25-μm line. Detailed balance calculations show that one expects a [S IX] 1.25-μm/3.75-μm intensity ratio of roughly 15 at log T_{e} (K) = 6 and log n_{e} (cm^{−3}) < 8. We find the following line flux estimates (× 10^{−18} W cm^{−2}): [S IX], 0.97 ± 0.08; Paβ, 11.45 ± 0.46.”

NOVA IN THE LARGE MAGELLANIC CLOUD 1992 Additional photometry (cf. IAUC 5669) by A. C. Gilmore obtained

with the 0.6-m f/16 Cassegrain at Mt. John University Observatory (reference star CPD −65^{o}433): Dec. 14.46 UT, V = 14.72 ± 0.02, U–B = –0.86 ± 0.11, B–V = –0.45 ± 0.05, V–R = +1.73 ± 0.04.

1992 December 28 Daniel W. E. Green